# Generalizing odd elasticity theory to odd thermoelasticity for planar materials

### 23 Jan 2024 v2

@@ I Introduction
The recent few years have witnessed the development of odd elasticity, a theoretical framework for elastic materials that do not store energy in the same way as hyperelastic materials do [1, 2, 3, 4] (for a review see [5]). Various physical systems display such responses, typically due to the breakdown of Maxwell-Betti reciprocity. At this point, new challenges arise. First, to extend the framework of odd elasticity to non-isothermal and/or non-adiabatic behaviors. This is considered here with three thermoelastic models: classical coupled thermoelasticity with Fourier law, thermoelasticity with relaxation times of Maxwell-Cattaneo type, and thermoelasticity with two relaxation times. While the latter two models are fundamentally different from one another, they can be simplified to the classical coupled one when the relaxation times are set to zero. The goal is to write down the governing equations admitting, in general, spatially inhomogeneous properties. The next question concerns the possibility of invariance of the stress field in an odd thermoelastic two-dimensional (planar) material. This problem of a so-called ”CLM shift in a compliance tensor field” [6] is tackled in a more general setting of eigenstrains. We determine an equivalence class of a wide range of odd materials in which the elastic compliances may be modified without affecting the stress field.
@@ II Irreversible thermodynamics
where $\varepsilon_{kl}=u_{(k},_{l)}$. This is analogous to viscoelasticity, where the stress or momentum current also contributes to entropy. Taking the divergence one gets
@@ III Thermoelasticity with parabolic or hyperbolic heat conduction
The Fourier model of heat conduction works very well when the signal propagation speed does not need to be accounted for; this is the basis of classical coupled (C-C) thermoelasticity. At sufficiently high modulation frequencies, the time delay between temperature gradient and heat flux needs to be considered, and this results in a Maxwell-Cattaneo model with one relaxation time, which is the basis of the Lord-Shulman (L-S) thermoelasticity, e.g. ([9, 10]). By contrast, the Green-Lindsay (G-L) thermoelasticity retains the Fourier law but accounts for two time delays in coupled field phenomena: one between the stress and stress-temperature tensor, and another between the entropy and the specific heat. Thus, with two relaxation times G-L offers a theory which is intermediate between C-C and L-S theories.
In the subsequent analysis, we focus on the thermal properties of odd elastic solids. This activity stems from breaking Maxwell-Betti reciprocity relations. Incorporation of thermal strains and stresses in elasticity leads to three basic models of thermoelasticity that generalize the above calassic models to the realm of odd solids:
@@
where, just like in (12), the overdot stands for the material time derivative [11]. As pointed out in that reference, this model is useful in various materials, typically polymeric, including the living [12] or dead [13] soft bio-tissues where $\tau$ may be on the order of one or several seconds. The resulting wave-type heat propagation is referred to as the second sound as opposed to the usual elastic waves (first sound). Also in this model, the Onsager symmetry may be broken.
@@
In the cases (i) and (ii), the thermoelastic constitutive law reads $\varepsilon_{ij}=S_{ijkl}\sigma_{kl}+A_{ij}\Delta T$, where $\varepsilon_{kl}=u_{(k},_{l)}$ is the strain with $\left(k,l\right)$ denoting symmetrization on indices $i$ and $j$, $u_{i}$ is the displacement; $S_{ijkl}$ is the compliance tensor (to be discussed below), and $\sigma_{kl}$ is the Cauchy stress tensor. Also, $A_{ij}$ ($=-S_{ijkl}M_{kl}$) is the thermal expansion tensor and $\Delta T$ is the temperature change from $T_{0}$. The thermoelasticity field equations can be compactly grasped in one system of two coupled equations for the displacement-temperature pair:
@@ V Stress field invariance under a shift in compliances
In the realm of materials science and mechanical engineering, the study of invariant properties of stress in plane elasticity and equivalence classes of composites plays a pivotal role. Plane elasticity simplifies stress and deformation to a two-dimensional perspective, essential for analyzing thin plates or surfaces under load. The invariants of stress, quantities unaltered by coordinate transformations, are crucial for comprehending material behavior under varying loads, aiding in predicting material failure and other mechanical properties. Concurrently, composites, engineered from multiple materials with distinct properties, are classified into equivalence classes based on shared characteristics such as mechanical or thermal behavior. This classification is vital for material selection in engineering, enabling the choice of composites that meet specific requirements like strength, weight, or heat resistance. Together, these concepts are fundamental in optimizing material properties for varied applications, ensuring safety and efficiency in high-stress environments or when employing advanced materials. Understanding stress invariance in the context of active odd solids can provide insights into their mechanical properties and stability, guiding their integration into engineering and material science applications where their unique characteristics might be advantageous.
@@
Inspecting (27) we see that, for the new stress field to remain $\sigma_{ij}$, there must hold these relations
@@
Put another way, although $S_{ijkl}\neq S_{klij}$, the answer to the above question is affirmative for a so-called shift of $S_{ijkl}$ to $\widehat{S}_{ijkl}$ according to
@@
$\left[\widehat{\mathbf{S}}\right]=\left[\mathbf{S}\right]+\left[\mathbf{S}^{I}\right],\text{ \ }\left[\mathbf{S}^{I}\right]=\frac{1}{2\Lambda}\left[\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&-2\end{array}\right].$ (32)
@@ VI Hydrodynamic modes
$\begin{split}v_{1}^{2}&=\frac{1}{\rho^{2}}[-(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)\rho\\ +&\sqrt{(8m_{1}m_{2}T_{0}\kappa_{o}+4m_{1}^{2}T_{0}\mu-4m_{2}^{2}T_{0}(B+\mu)+(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)^{2}+4c_{e}(\kappa_{o}^{2}+\mu(B+\mu)))\rho}].\end{split}$ (34a)
$\begin{split}v_{2}^{2}&=\frac{1}{\rho^{2}}[-(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)\rho\\ -&\sqrt{(8m_{1}m_{2}T_{0}\kappa_{o}+4m_{1}^{2}T_{0}\mu-4m_{2}^{2}T_{0}(B+\mu)+(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)^{2}+4c_{e}(\kappa_{o}^{2}+\mu(B+\mu)))\rho}].\end{split}$ (34b)
@@
$\Gamma_{1}=-2\frac{2k_{1}(4\kappa_{o}^{2}+(2\mu-v_{1}^{2}\rho)(2B+2\mu-v_{1}^{2}\rho))}{v_{1}^{2}(v_{1}^{2}-v_{2}^{2})\rho^{2}},$ (35a)
$\Gamma_{2}=-2\frac{2k_{1}(4\kappa_{o}^{2}+(2\mu-v_{2}^{2}\rho)(2B+2\mu-v_{2}^{2}\rho))}{v_{2}^{2}(v_{2}^{2}-v_{1}^{2})\rho^{2}},$ (35b)
@@ Acknowledgments
M.O.-S. thanks the Isaac Newton Institute for Mathematical Sciences, University of Cambridge, for support and hospitality as the Rothschild Distinguished Visiting Fellow during the programme ”Uncertainty quantification and stochastic modelling of materials” where work on this paper was partially completed. P.S. acknowledges support from the Polish National Science Centre (NCN) Sonata Bis grant 2019/34/E/ST3/00405. Part of this work was performed at the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452.

### 22 Feb 2024 v3

@@ I Introduction
The recent few years have witnessed the development of odd elasticity, a theoretical framework for elastic materials that do not store energy in the same way as hyperelastic materials do [1, 2, 3, 4] (for a review see [5]). Various physical systems display such responses, typically due to the breakdown of Maxwell-Betti reciprocity. At this point, new challenges arise. First, to extend the framework of odd elasticity to non-isothermal and/or non-adiabatic behaviors. This is considered here with three thermoelastic models: classical coupled thermoelasticity with Fourier law, thermoelasticity with relaxation times of Maxwell-Cattaneo type, and thermoelasticity with two relaxation times. While the latter two models are fundamentally different from one another, they can be simplified to the classical coupled one when the relaxation times are set to zero. The goal is to write down the governing equations admitting, in general, spatially inhomogeneous properties. The next question concerns the possibility of invariance of the stress field in an odd thermoelastic two-dimensional (planar) material. This problem of a so-called ”Cherkaev-Lurie-Milton (CLM) shift in a compliance tensor field” [6] is tackled in a more general setting of eigenstrains. We determine an equivalence class of a wide range of odd materials in which the elastic compliances may be modified without affecting the stress field.
@@ II Irreversible thermodynamics
where $\varepsilon_{kl}=u_{(k,l)}$. This is analogous to viscoelasticity, where the stress or momentum current also contributes to entropy. Taking the divergence one gets
@@ III Thermoelasticity with parabolic or hyperbolic heat conduction
The Fourier model of heat conduction works very well when the signal propagation speed does not need to be accounted for; this is the basis of classical coupled (C-C) thermoelasticity. At sufficiently high modulation frequencies, the time delay between the temperature gradient and heat flux needs to be considered, and this results in a Maxwell-Cattaneo model with one relaxation time, which is the basis of the Lord-Shulman (L-S) thermoelasticity, e.g. [9, 10]. By contrast, the Green-Lindsay (G-L) thermoelasticity retains the Fourier law but accounts for two time delays in coupled field phenomena: one between the stress and stress-temperature tensor, and another between the entropy and the specific heat. Thus, with two relaxation times G-L offers a theory which is intermediate between C-C and L-S theories.
In the subsequent analysis, we focus on the thermal properties of odd elastic solids. This activity stems from breaking Maxwell-Betti reciprocity relations. Incorporation of thermal strains and stresses in elasticity leads to three basic models of thermoelasticity that generalize the above classic models to the realm of odd solids:
@@
where, just as in (12), the overdot stands for the material time derivative [11]. As pointed out in that reference, this model is useful in various materials, typically polymeric, including the living [12] or dead [13] soft bio-tissues where $\tau$ may be on the order of one or several seconds. The resulting wave-type heat propagation is referred to as the second sound as opposed to the usual elastic waves (first sound). Also in this model, the Onsager symmetry may be broken. Hereinafter, we switch to the standard tensor notation for a gradient: a comma followed by the index of coordinate.
@@
In cases (i) and (ii), the thermoelastic constitutive law reads $\varepsilon_{ij}=S_{ijkl}\sigma_{kl}+A_{ij}\Delta T$, where $\varepsilon_{kl}=u_{(k},_{l)}$ is the strain with $\left(k,l\right)$ denoting symmetrization on indices $i$ and $j$, $u_{i}$ is the displacement; $S_{ijkl}$ is the compliance tensor (to be discussed below), and $\sigma_{kl}$ is the Cauchy stress tensor. Also, $A_{ij}$ ($=-S_{ijkl}M_{kl}$) is the thermal expansion tensor and $\Delta T$ is the temperature change from $T_{0}$. The thermoelasticity field equations can be compactly grasped in one system of two coupled equations for the displacement-temperature pair:
@@ V Stress field invariance under a shift in compliances
In the realm of materials science and mechanical engineering, the study of invariant properties of stress in plane elasticity and equivalence classes of composites plays a pivotal role. Plane elasticity simplifies stress and deformation to a two-dimensional perspective, essential for analyzing thin plates or surfaces under load. The invariants of stress, quantities unaltered by coordinate transformations, are crucial for comprehending material behavior under varying loads, aiding in predicting material failure and other mechanical properties. Concurrently, composites, engineered from multiple materials with distinct properties, are classified into equivalence classes based on shared characteristics such as mechanical or thermal behavior. This classification is vital for material selection in engineering, enabling the choice of composites that meet specific requirements such as strength, weight, or heat resistance. Together, these concepts are fundamental in optimizing material properties for varied applications, ensuring safety and efficiency in high-stress environments or when employing advanced materials. Understanding stress invariance in the context of active odd solids can provide insights into their mechanical properties and stability, guiding their integration into engineering and material science applications where their unique characteristics might be advantageous.
@@
Inspecting (27) we see that, for the new stress field to remain $\sigma_{ij}$, these relations must hold
@@
In other words, although $S_{ijkl}\neq S_{klij}$, the answer to the above question is affirmative for a so-called shift of $S_{ijkl}$ to $\widehat{S}_{ijkl}$ according to
@@
$\left[\widehat{\mathbf{S}}\right]=\left[\mathbf{S}\right]+\left[\mathbf{S}^{I}\right],\text{ \ }\left[\mathbf{S}^{I}\right]=\frac{1}{2\Lambda}\begin{bmatrix}0&1&0\\ 1&0&0\\ 0&0&-2\end{bmatrix}$ (32)
@@ VI Hydrodynamic modes
$\begin{split}v_{1}^{2}&=\frac{1}{\rho^{2}}[-(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)\rho\\ +&\sqrt{(8m_{1}m_{2}T_{0}K^{0}+4m_{1}^{2}T_{0}\mu-4m_{2}^{2}T_{0}(B+\mu)+(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)^{2}+4c_{e}((K^{0})^{2}+\mu(B+\mu)))\rho^{2}}].\end{split}$ (34a)
$\begin{split}v_{2}^{2}&=\frac{1}{\rho^{2}}[-(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)\rho\\ -&\sqrt{(8m_{1}m_{2}T_{0}K^{0}+4m_{1}^{2}T_{0}\mu-4m_{2}^{2}T_{0}(B+\mu)+(Bc_{e}+(m_{1}^{2}-m_{2}^{2})T_{0}+2c_{e}\mu)^{2}+4c_{e}((K^{0})^{2}+\mu(B+\mu)))\rho^{2}}].\end{split}$ (34b)
@@
$\Gamma_{1}=-2\frac{2k_{1}(4(K^{0})^{2}+(2\mu-v_{1}^{2}\rho)(2B+2\mu-v_{1}^{2}\rho))}{v_{1}^{2}(v_{1}^{2}-v_{2}^{2})\rho^{2}},$ (35a)
$\Gamma_{2}=-2\frac{2k_{1}(4(K^{0})^{2}+(2\mu-v_{2}^{2}\rho)(2B+2\mu-v_{2}^{2}\rho))}{v_{2}^{2}(v_{2}^{2}-v_{1}^{2})\rho^{2}},$ (35b)
@@ Acknowledgments
M.O.-S. thanks the Isaac Newton Institute for Mathematical Sciences, University of Cambridge, for support and hospitality as the Rothschild Distinguished Visiting Fellow during the program ”Uncertainty quantification and stochastic modelling of materials” where work on this paper was partially completed. P.S. acknowledges support from the Polish National Science Centre (NCN) Sonata Bis grant 2019/34/E/ST3/00405. Part of this work was performed at the Aspen Center for Physics, which is supported by the National Science Foundation grant PHY-2210452.